40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16268     Entry time: Tue Aug 3 20:20:08 2021
Author: Anchal 
Type: Update 
Category: Optical Levers 
Subject: Recentered ETMX, ITMX and ETMY oplevs at good state 

Late elog. Original time 08/02/2021 21:00.

I locked both arms and ran ASS to reach to optimum alignment. ETMY PIT > 10urad, ITMX P > 10urad and ETMX P < -10urad. Everything else was ok absolute value less than 10urad. I recentered these three.

Than I locked PRMI, ran ASS on PRCL and MICH and checked BS and PRM alignment. They were also less than absolute value 10urad.

ELOG V3.1.3-